if f(x)=X^2-6x + 2 and g(x)=-2x, find the following composition.
(g f)(5)

Answer by amarjeeth123(574)

gof(x)=g(f(x))=g(x^2-6x + 2)=-2(x^2-6x + 2)
gof(5)=-2(25-30+2)=-2(-3)=6